# Let’s Get to the Point: LLM-Supported Planning, Drafting, and Revising of Research-Paper Blog Posts

### v1

@@ 2.4. Scientific Blog Posts
Scientists can engage with broad community through various channels, including press releases, magazines, journals, and tweets (August et al., 2020; Gero et al., 2022). This work focuses on scientific blog posts, which offers increased recognition and transparency of one’s work, writing practice, editorial freedom, and public engagement (Jarreau, 2015; Blanchard, 2011; Zou and Hyland, 2019). There are several forms of scientific blog posts, including academic commentary, free commentary, political commentary, and mediation of research to laypersons (Mahrt and Puschmann, 2014). Here, we focus on blog posts with the primary goal of communicating information from a research paper to scientists in the broad area of the paper (e.g., computer science, psychology).
3. Formative Study and System Iteration
We conducted a formative study to identify challenges that people encounter in working with an LLM-powered technology probe for writing blog posts about their research papers, and corresponding design goals for our final LLM tool.
3.1. Participants
Via academic social networks, we recruited 6 participants for the formative study and compensated them with $25 for the 45-minute study. Participants were PhD students, a population frequently tasked with writing blog posts for papers they have authored. Five worked in computer science and one in information science. To circumvent initial time required for getting used to LLMs themselves, we recruited people who already had some experience of working with LLMs in general (6-10 sittings: 1; ¿10 sittings: 5). We also required participants to have previously written research papers (3-10 papers: 3; ¿10 papers: 3), since the purpose of such a tool is to write blog posts for one’s own research papers. Participants had a mixture of experience with writing research-paper blog posts (0 posts: 4; 1-2 posts: 2). Also, for sufficient participant motivation, we made sure that all participants were more interested than not (at least 5 on a 7-point Likert scale where 7 indicates the highest interest) in writing blog posts for their research papers.
3.2. Study Procedure
The study consisted of 3 iterations, each of which involved sessions with 2 participants. Based on participants’ interactions and feedback, we updated the probe after each iteration for better insight into remaining challenges. Prior to each study session, the participant completed a participation agreement. The session coordinator recorded and transcribed each session in Google Hangouts. The coordinator asked the participant to assume their audience was laypeople, as a common advantage of writing scientific blog posts is to communicate with a broader audience than the paper does (Jarreau, 2015; Blanchard, 2011). The participant then used a link to access the technology probe initialized for one of their research papers that they provided. In order to see how participants would intuitively use the probe, the coordinator purposefully did not initially explain how to use it. For 30 minutes, the participant used the probe to write as much as possible for a blog post about their research paper. We asked participants to think aloud as they worked (Van Someren et al., 1994). The coordinator occasionally asked follow-up questions about what the participant said aloud. Participants could ask for help when they did not understand an aspect of the tool. Once 30 minutes had passed, the coordinator engaged the participant in a semi-structured interview. They asked the participant about the difficulties and benefits of using the probe to 1) navigate the paper and outline, 2) generate and edit text for a blog post section, and 3) edit the blog post layout. Finally, they asked for the participant’s thoughts on receiving different types of feedback on their blog post, but this topic was left for future work.
3.3. Initial Technology Probe
In the first iteration of the technology probe interface (Appendix 28), the left side consisted of blog post sections pre-populated with an LLM-generated draft for drafting and revising the blog post. The right side contained the paper text with selectable paragraphs and an LLM-generated paper outline with bullet points for planning. Under the currently selected blog post section, there was an LLM-supported text generator that contained buttons to generate section text from the selected paragraphs and bullet points as well as to expand and condense the text in the blog post section editor.
3.4. Design Goals and Tool Iteration
After each iteration of the formative study, the first author conducted inductive thematic analysis (Braun and Clarke, 2006) of the participants’ think-aloud and interview comments on their experience with the probe. Participants described difficulties in working with the probe that, aside from general usability issues, may be classified under two overarching challenges.
3.4.1. Challenge 1: Planning Paper Content to Use
An academic paper is a long document with dense information, so we expected that identifying information to include in the blog post would be a challenge. For this reason, our technology probe parsed the paper text and generated a paper outline with an LLM in order to support navigating and utilizing the paper. Still, participants faced difficulties with finding and selecting information from the paper. Our associated design goal was to help writers plan their writing by facilitating review of a long document’s contents and selection of essential information to include in the writing (DG1).
Although participants faced issues with the paper outline, four participants expressed that they liked having the paper outline, and four appreciated the ability to select paper paragraphs or bullet points from which to generate text. Because participants overall liked and used the paper outline for identifying and using essential information from the paper, we iterated on this affordance to address DG1. We identified the following requested functionality associated with this design goal:
Transparent Connection Between Paragraphs and Bullet Points. In the first iteration, P2 noted confusion around how the bullet points in the outline corresponded to the original paper text, despite the existence of a button to the left of each set of bullet points that revealed the associated paragraph when clicked. We therefore added a bright line to the left of each set of bullet points to better visualize its grouping and its connection to the corresponding paragraph (see Figure 2). In the second iteration, P3 expressed uncertainty about how the selection of paragraphs and bullet points were related. For example, they were surprised that a paragraph was selected but only one of three corresponding bullet points was selected. During the formative study, the selection logic was as follows. If a bullet point is selected, the associated paragraph needs to be selected, but if a paragraph is selected, no bullet point needs to be selected. This was done in an attempt to make it clearer that the associated paragraph was given to the model as context in addition to any bullet points selected. However, this logic seemed to confuse the participants.
Filter for Selected Content. In the third iteration, P6 said that they were not sure if they had selected everything that they wanted to select. Throughout all the iterations, the probe only communicated the number of selected paragraphs and bullet points. P6 desired a way to view only the selected paragraphs and bullet points.
3.4.2. Challenge 2: Drafting and Revising with Recurring LLM Instructions
The second challenge was attempting to provide the LLM instructions for generating and modifying text that were not well-supported by the probe. Some of these instructions were frequently desired, but the participant had no reasonable way to input them or, once directly prompting the LLM was available, had to determine a prompt to fit their needs. Participants also faced confusion regarding the prompts being used in the back-end. Our corresponding design goal was therefore to help writers draft and revise their writing by transparently streamlining commonly desired instructions to the LLM for generating and modifying text (DG2). We identified the following requested functionalities under this design goal:
Transparent Prompt Input. In the second iteration, P3 liked a default blog post section and was confused that it appeared significantly different from the new version they generated without adjusting the pre-selected bullet points. For one, there was a difference in the requested length in the prompt for the initial-draft section versus the live-generation section. Also, the participant was not aware of the section-specific instructions that were provided to the model for the initial-draft section. To provide access to default settings, which may often be desirable, we updated the probe to show these section-specific instructions as preset (but adjustable) custom instructions. While the model was always prompted to add 1 to 2 sentences in the third iteration, P5 was confused regarding if and how much the text would expand after clicking the ”expand” button multiple times. Meanwhile, P6 commented that they were unsure if the model had context for the blog post sections above the one for which they were generating text.
Custom Instructions. In the first iteration, our design probe did not support direct custom instructions. Users could only generate or modify the text using the pre-defined operations. P1 tried to work around this limitation by giving specific instructions to the LLM in the blog post section editor, custom bullet points, and custom instructions. The participant seemed unaware that the model was not taking these instructions into account. To provide users with more control over the generated text, we added a field for custom instructions to the generator box.
New Section Generation. In the second iteration, P4 noted wanting to generate entire new sections with the model, perhaps based on the section header.
Flexible Input for Text Modification. In the first iteration, P2 expected and wanted the ”condense” button to iterate on the generated text rather than the editor text, so we updated the ”expand” and ”condense” buttons to do that. However, in later iterations, P4 expressed interest in the model expanding on something they wrote without reference to the paper, and P6 expressed interest in rewriting portions of the original paper in different manners.
Preset Modification Actions. We observed participants attempting to rewrite text in manners that seemed applicable to many research-paper blog posts. For instance, P1 wanted a button to make text more approachable for their audience of laypeople, while P3 rewrote text that seemed ”too lofty” in its claims. For the third iteration, we therefore added buttons for such modifying actions, which included ”simpler terms”, ”less dramatic”, and ”more dramatic.” When the participant clicked the button, the modification would be applied to the last generation.
@@ 4. The Papers-to-Posts System
4.1. Paper Outline and Initial Draft Generation
@@
The paper text was extracted from the paper PDF using a public tool for converting scientific PDFs to HTML.111https://papertohtml.org We manually adjusted the parsed HTML so that the section headers were correct, no section or subsection was completely missing, and only the main paper text was included (i.e., no appendix, references, or footnotes). Within the paper text box, participants could click a button to see a paper outline generated for the paper text. To obtain this paper outline, we used an LLM to generate 1 to 3 bullet points to summarize each paragraph in the paper, depending on the length of the paragraph. To generate the initial blog post draft for the tool, all of the paper’s bullet points were given to the model along with the request to select 10 bullet points relevant to the initial four blog post sections: introduction, methods, results, and conclusion. To generate each draft section, the model was provided the selected bullet points, their associated paragraphs, the portion of the blog post that had already been generated, and section-specific guidelines based on prior work (Mehlenbacher, 2019), popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]; Medium, [n. d.]), and feedback from three communications experts. We found that this process of generating one draft section at a time with the context of prior sections helped make the draft more coherent while avoiding overwhelming the model with instructions for many sections. For both the initial draft and user generations, the model was provided the associated paragraphs for any bullet points in order to ground the model in the actual paper. We used GPT-4 for all LLM tasks except for extracting bullet points from a paragraph, which we produced using GPT-3.5-Turbo after observing that it produced shorter bullet points that were quicker to read. It was also many times cheaper for this token-intensive generation task.
4.2. User Interface
Developed over the course of the formative study, the Papers-to-Posts system consists of two main panes split vertically. On the left is the pane for writing the blog post (Figure 2a), and on the right is the area for generating and modifying text. The blog post is divided into sections, which can be named, reorganized, and deleted. Each section is connected to a unique LLM state, so the history of all inputs and outputs for text generation and modification are stored separately for each section. The initial blog post draft is generated as described in Section 3.3. The blog post section currently being edited is highlighted in purple to indicate that it is the section for which text is generated. Using the controls on the top of the left pane, users can view the blog post’s word count, a preview of the full blog post, and tips for writing blog posts based on prior work (Mehlenbacher, 2019) and popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]; Medium, [n. d.]), as well as feedback from 3 communications experts. There is a button to submit the blog post at the bottom. The writer can also add a new section to the blog post using the plus-sign button next to each blog post section. When the writer clicks this button, they can either create a blank section or generate a section based on the section header they provide. To generate a new section, we prompt GPT-4 to select 10 bullet points for the section based on the provided header and to generate text for the section focused on those bullet points (Appendix A).
@@ 4.2. User Interface
The pane for generating and modifying text on the right has three areas. The first is the planning-support area (Figure 2b), which contains the Paper Outline to address DG1. The Paper Outline supports writers in planning their blog post, through review of the paper’s contents and selection of essential information to include in the post. The paper view at the top of the pane offers the writer two views of the paper in two tabs: Paper Text and Paper Outline. Under the Paper Text tab, the writer may view the original paper text, divided into paragraphs. The paper is parsed into collapsible sections. The figures and tables are presented in a separate final section of the paper. Their captions can be copied for pasting in the editor so that the writer knows where they should go in the final blog post. Under the Paper Outline tab, the writer may view the paper outline, divided into sets of bullet points associated with the original text’s paragraphs. The writer may view a bullet point set’s corresponding paragraph by clicking the button to the left, and the reverse is true under the Paper Text tab. The outline is generated, and the pre-selected bullet points for each blog post section are displayed. The writer may select or deselect any bullet points or paragraphs that they want the LLM to use in generating text for the blog post section on which they are working. Selecting a paragraph does not affect the selection of associated bullet points, and vice versa. On the back-end, the model still receives the context of the associated paragraph for each bullet point selected. There are buttons below the paper view to clear all selected bullet points and paragraphs, filter for only selected bullet points and paragraphs, and close or open all the paper sections. When the filter button is activated, if the writer is in the Paper Text tab, if a paragraph was deselected but a corresponding bullet point is still selected, the tool notes that ”A corresponding bulletpoint(s) is still selected,” and vice versa.
Below the planning-support area is the drafting-support area (Figure 2c). There are several inputs to address DG2 by transparently streamlining commonly desired instructions to the LLM for generating text. First, there is a menu to select the desired length for the generated text. The default length is labeled as ”auto,” and the other options are 1 sentence, 1 paragraph, or a few paragraphs. The ”auto” length is the same as for the initial LLM draft sections (around 125 to 250 words) There is next a toggle turned on by default for the LLM to use the context of blog post sections above the one currently being edited. The writer can optionally utilize text fields to provide custom bullet points, custom instructions, and starting text to the LLM. For the initial introduction, methods, results, and conclusion sections of the blog post, the custom instructions field is pre-filled with the section-specific instructions used to generate those sections. The text generator’s ”generate” button utilizes the inputs to generate text for the current blog post section using GPT-4.
The tool also provides a revising-support area (Figure 3d, located below the drafting-support area in the tool), which addresses DG2 by transparently streamlining commonly desired instructions to the LLM for modifying text. First, there is a text field for inputting the text to be modified. This can be anything, including text from the blog post draft, the paper, or a generation. There is then a menu to select the desired length for the modified text. The options are the same as for generated text, with ”auto” being unique to each modifying action (see Appendix A). Next, there is a toggle, on by default, for the LLM to use the context of blog post sections above the one currently being edited. There is also a toggle turned on by default for the LLM to base modified writing on selected paragraphs and bullet points. Underneath the toggles, there is a text field for providing any custom instructions for the modification action. At the bottom, there are 5 preset modification buttons: ”expand”, ”condense”, ”simpler terms”, ”less dramatic”, and ”more dramatic”. There is also a ”custom modification” button. Aside from the ”expand” and ”condense” buttons, all of the affordances for generating and modifying text were additions made to support recurring types of LLM instructions observed in the formative study. In the lab study, participants utilized each affordance to some degree.
5. Lab Study
@@
We conducted a lab study to determine if and how the LLM tool resulting from the formative study may support researchers in writing blog posts about their research papers.
5.1. Hypotheses

### v2

@@ 2.5. Scientific Blog Posts
Scientists can engage with broad community through various channels, including press releases, magazines, journals, and tweets (August et al., 2020; Gero et al., 2022). This work focuses on scientific blog posts, which offer increased recognition and transparency of one’s work, writing practice, editorial freedom, and public engagement (Jarreau, 2015; Blanchard, 2011; Zou and Hyland, 2019). There are several forms of scientific blog posts, such as academic commentary, free commentary, and mediation of research to laypersons (Mahrt and Puschmann, 2014). Here, we focus on blog posts with the primary goal of communicating information from a research paper to people unfamiliar with the paper’s specific research topic.
@@ 3. The Papers-to-Posts System
We developed Papers-to-Posts based on key design goals and pilot study feedback, which helped us to identify system features to add for increased transparency, flexibility, and ease of use. This section describes the design goals for Papers-to-Posts and provides a walkthrough of the resulting system’s four main areas.111A walkthrough video of the tool may be found in the supplementary materials.
3.1. Design Goals
We determined the design goals for Papers-to-Posts based on guidelines for academic blog posts from popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]), prior work (Mehlenbacher, 2019), and feedback from three communications experts, recruited from our institutions. There are multiple guidelines on the content and format for different sections of the academic blog post (e.g., introduction, results) (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]; Mehlenbacher, 2019). With respect to the blog post’s tone, guidelines suggest avoiding jargon (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]), sensationalism (noted by a communications expert), and wordiness (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]).
The resulting design goals are:
DG1: Help researchers to plan what paper content to include in their blog post.
DG2: Help researchers to draft their content in sections amenable to a research-paper blog post.
DG3: Help researchers to revise the blog post’s tone to be scientifically accurate but also concise, easy to understand, and engaging.
3.2. Walkthrough of Start-Up Step
3.2.1. Frontend
In the Blog Post Area (Figure 2a), the user starts by reviewing the initial blog post draft. The blog post draft is divided into sections. The user clicks on whichever section they want to work on first, and the tool highlights that section for reference. At the top, the user views the blog post’s word count, a preview of the full blog post, and tips for writing blog posts, which are based on popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]), prior work (Mehlenbacher, 2019), and feedback from three communications experts.
3.2.2. Backend
@@
In order to obtain the initial blog post draft for a given research paper, we completed a few steps (Figure 1a). In the first step, we used a public tool for converting scientific PDFs to HTML222https://papertohtml.org to extract the paper text from the paper PDF. For the lab study, we then manually adjusted the parsed HTML so that the section headers were correct, no section or subsection was completely missing, and only the main paper text was included (i.e., no appendix, references, or footnotes). We also manually retrieved the tables and figures from the parsing. For the real-world deployment study, the parsing was automatic to better align with a real-world setting. For the second step, we used GPT-3.5 Turbo to generate between one and three bullet points to summarize each paragraph in the parsed paper, depending on the length of the paragraph. While we used GPT-4 (GPT-4-0125-preview in the deployment study) for all of Papers-to-Posts’s other tasks, we decided to use GPT-3.5-Turbo for this task for two reasons: 1) it produced shorter bullet points that were quicker to read, and 2) it was many times cheaper for this token-intensive generation task. All together, the generated bullet points formed a paper outline. For the third step, we provided the LLM with all of the paper’s bullet points along with the request to select 10 bullet points relevant to the initial four blog post sections: introduction, methods, results, and conclusion. We requested 10 bullet points per blog post section in order to provide the model sufficient content without overwhelming it. For the fourth step, we prompted the LLM to generate each section in succession using the selected bullet points for that section, the paragraphs associated with the selected bullet points (to better ground the model), the portion of the blog post that had already been generated, and section-specific guidelines based on popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]), prior work (Mehlenbacher, 2019), and feedback from three communications experts. We observed that this process of generating one draft section at a time with the context of prior sections helped to make the draft more coherent while avoiding overwhelming the model with instructions for multiple sections. The model was prompted to make each section around 125 to 250 words and around one to three paragraphs. The prompts for generating the initial draft may be found in Appendix B.2. An example initial draft and associated selected bullet points is in Appendix F.1 and F.2.
3.3. Walkthrough of Planning Step
3.3.1. Frontend
@@
In the Planning-Support Area (Figure 2b), the user works with the tool to plan the blog post section on which they are currently working (DG1). This area has two tabs labeled Paper Outline and Paper Text. Under the Paper Outline tab, the user views the paper’s outline, divided into sets of one to three bullet points; each set is associated with a paragraph from the original text. The user views a bullet point set’s corresponding paragraph by clicking the button to the left, and the reverse is true under the Paper Text tab, where the user views the original paper text divided into paragraphs. Under both tabs, the information is divided into collapsible sections based on the paper’s highest-level sections.333In the lab study, the tool presented the figures and tables in a final collapsible section. Their captions could be copied for pasting in the editor to help the user plan where to put figures in the final blog post. We did not include this feature in the deployment study, as the automatic paper parsing did not extract this information.
Under the Paper Outline tab, the user reviews the tool’s pre-selected bullet points for the current blog post section by scrolling through the full outline or clicking the “Show Selected Only” button below the paper area. When this filter button is activated and the user is under the Paper Text tab, if a paragraph is not selected but a corresponding bullet point is selected, the tool notes that “A corresponding bulletpoint(s) is still selected”; the opposite is true for the Paper Outline tab. The user clicks the “Close All Paper Sections” button below the paper area if they want to get a higher-level view of the paper. The user then adjusts the selection of bullet points or paragraphs for the LLM to use in generating text for the current blog post section. Selecting a paragraph does not affect the selection of associated bullet points, and vice versa. If the user wants to reset the content selection, they click the “Clear All Selected” button below the paper area.
3.3.2. Backend
As shown in Figure 1b, all planning actions (as well as drafting and revising actions) are connected to the user’s selected blog post section. In other words, the state and history of all inputs and outputs for text generation are stored separately for each section. This is somewhat similar to the fragmented document history proposed by Buschek (Buschek, 2024).
3.4. Walkthrough of Drafting Step
3.4.1. Frontend
In the Drafting-Support Area (Figure 2c), the user works with the tool to synthesize text for the blog post section on which they are currently working (DG2). The user reviews and adjusts the LLM’s section-specific instructions for generating text. At the top of the Drafting Support Area, the user changes the desired length of the generated text to their liking. The length defaults to “auto” and may be updated to one sentence, one paragraph, or a few paragraphs. Below, the user notes the toggle that is turned on by default. This toggle lets the user know that the context of blog post sections above the highlighted one is provided to the LLM in order to generate text for this section, and they may turn this toggle off if they do not want this context to influence the generated text for this section. In the “Custom Bulletpoints” text field, the user enters any additional bullet points that they want their generated text for the highlighted section to cover. In the “Custom Instructions” text field, the user adjusts the instructions for the highlighted section’s generated text. For each of the initial blog post sections, this box is pre-filled with section-specific instructions, which are based on the blog post guidelines we collected. In the text field labelled “Starting Text to Continue from”, the user optionally provides text that should be at the start of the generated text. Finally, the user clicks the “generate” button to generate text for the highlighted section, which appears below. The user then copies and pastes the generated text into the highlighted blog post section and edits it as they see fit. The user may peruse a history of their content selection, LLM instructions, and text generations for the highlighted section.
The user also drafts text for the blog post by adding a new section. In the Blog Post Area, the user clicks the plus-sign button next to an existing blog post section, and the tool places the new section below the existing one. The tool gives the user the option to either create a blank section or generate a section based on the section header that they provide. If the user decides the new section should be in another location within the blog post, they use the section’s associated up and down arrow buttons to move the section to the desired point. For this new section, the user may return to the planning step or adjust the LLM’s instructions for generating text. If the user decides not to include the section in the blog post, they delete it using the trash button.
3.4.2. Backend
As shown in Figure 1c, the LLM (GPT-4 in lab study; GPT-4-0125-preview in deployment) generates text for the highlighted blog post section with the user’s adjusted selection of bullet points and paragraphs as well as their adjusted section-specific instructions (described in Section 3.4.1). With respect to the selected bullet points, the model receives the context of the associated paragraph for each bullet point selected, even if the associated paragraph was not selected by the user. This is to ensure the model’s generation is grounded in the actual paper text. When the user desires the “auto” length for generated text, the model is instructed to generate text that is 125 to 250 words and between one and three paragraphs. To generate a new section, we prompt the LLM to select 10 bullet points for the section based on the provided header and to generate text focused on those bullet points. Again, the model is provided the context of the bullet points’ associated paragraphs. The prompts for drafting text may be found in Appendix B.4, and example outputs are in Appendix F.1 and F.2.
3.5. Walkthrough of Revising Step
3.5.1. Frontend
In the Revising-Support Area (Figure 2d), the user works with the tool to revise the blog post section on which they are currently working (DG3). At the top of this area, the user pastes the text that they want to modify in the “Writing to Modify” text field. This text may be anything, including text from the blog post draft, the paper, or a previous model generation. The user then adjusts the desired length of the modified text if need be; the options are the same as in the Drafting-Support Area (Section 3.4). Also as in the Drafting-Support Area, the user may adjust the first toggle, which is on by default to indicate that the blog post sections above the one being edited are being provided as context to the LLM. The user then notes the second toggle, which is on by default to indicate that the LLM will utilize the currently selected paragraphs and bullet points to generate the modified text. At the bottom of the Revising-Support Area, there are five preset modification buttons to 1) expand the text, 2) condense the text, 3) rewrite the text in simpler terms to be more understandable to a layperson, 4) rewrite the text in a less dramatic tone to align better with the unadorned language of scientific writing, and 5) rewrite the text in a more dramatic tone to better capture readers’ attention. There is also a “custom modification” button. If the user wants to use the “custom modification” button, they enter their custom instructions for the text modification in the text field above. If the user wants to use a different modification button, they may still provide custom instructions for the LLM to take into consideration. The user copies and pastes the modified text into the highlighted blog post section and edits it as they see fit. The user may peruse a history of their LLM instructions and text modifications for the highlighted section.
3.5.2. Backend
As shown in Figure 1d, the LLM (GPT-4 in lab study; GPT-4-0125-preview in deployment) modifies text for the highlighted blog post section with the user’s selected text snippet and their adjusted instructions for the modification (described in Section 3.5.1), including the selected text modification. The desired length “auto” is unique to each modifying action. The prompts for revising text may be found in Appendix B.4, and example outputs are in Appendix F.3.
4. Evaluation Studies
4.1. Lab Study
@@
We conducted a lab study to determine if and how Papers-to-Posts may support researchers in writing blog posts about their research papers.
4.1.1. Hypotheses